40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 6 01:09:11 2017, johannes, Update, Computer Scripts / Programs, New set of loss measurements XARM_20170905.pdfYARM_20170905.pdf
    Reply  Mon Sep 11 12:56:40 2017, johannes, Update, Computer Scripts / Programs, lossmap attempts 
       Reply  Mon Sep 11 16:46:09 2017, Steve, Update, PEM, earthquakes M5.4eq.png
          Reply  Tue Sep 19 14:00:41 2017, Steve, Update, PEM, earthquakes 7.1MX5.3J.png
             Reply  Fri Oct 27 09:34:20 2017, Steve, Update, PEM, earthquakes recentEQs.png
                Reply  Fri Nov 10 13:46:26 2017, Steve, Update, PEM, M3.1 local earthquake 3.1M_local_EQ.png
                   Reply  Mon Apr 9 08:39:39 2018, Steve, Update, PEM, M5.3 eq Souther CA M5.3_Santa_Cruz_Is.CA.pngafter_M5.3.pngM5.3vac.png
                Reply  Thu Jan 25 08:27:25 2018, Steve, Update, PEM, earthquakes M4_local_eq.pngM4.pngM5.8Ferndale_CA.png
       Reply  Mon Sep 25 17:56:33 2017, johannes, Update, Computer Scripts / Programs, transmitted power during lossmap 
    Reply  Thu Oct 11 13:40:51 2018, yuki, Update, Computer Scripts / Programs, loss measurements 
       Reply  Fri Oct 12 12:29:34 2018, yuki, Update, Computer Scripts / Programs, loss measurements 
          Reply  Fri Oct 12 20:20:29 2018, yuki, Update, Computer Scripts / Programs, loss measurements 
             Reply  Sat Oct 13 20:11:10 2018, yuki, Update, Computer Scripts / Programs, loss measurements 
                Reply  Mon Oct 15 10:32:13 2018, yuki, Update, Computer Scripts / Programs, loss measurements XARM_20181015_1500.pdfYARM_20181015_1500.pdf
                   Reply  Tue Oct 16 00:44:29 2018, yuki, Update, Computer Scripts / Programs, loss measurements 
                   Reply  Fri Nov 2 19:25:16 2018, gautam, Update, Computer Scripts / Programs, loss measurements 
    Reply  Mon Oct 15 12:52:54 2018, yuki, Update, Computer Scripts / Programs, additional comments 
       Reply  Fri Nov 2 16:42:31 2018, aaron, Update, Computer Scripts / Programs, arm loss measuremenents 
          Reply  Mon Nov 5 13:52:18 2018, aaron, Update, Computer Scripts / Programs, arm loss measuremenents 
             Reply  Tue Nov 6 10:19:26 2018, aaron, Update, Computer Scripts / Programs, arm loss measuremenents 
                Reply  Tue Nov 6 15:23:48 2018, gautam, Update, IOO, IMC problematic MCwonky.png
                   Reply  Tue Nov 6 19:02:35 2018, aaron, Update, IOO, IMC problematic 181106_MCTRANS.jpg
                      Reply  Sat Nov 10 17:40:00 2018, aaron, Update, IOO, IMC problematic 
                         Reply  Mon Nov 12 13:53:20 2018, rana, Update, IOO, loss measurement: oscope vs CDS DAQ 
                         Reply  Thu Nov 15 10:21:07 2018, aaron, Update, IOO, IMC problematic 181115_chansDown.pngPD_noise.png
                            Reply  Fri Nov 16 10:53:07 2018, aaron, Update, IOO, IMC problematic Screenshot_from_2018-11-16_19-22-34.png
                               Reply  Sat Nov 17 18:59:01 2018, aaron, Update, IOO, IMC problematic 
                Reply  Wed Nov 7 05:16:16 2018, yuki, Update, Computer Scripts / Programs, arm loss measuremenents 
Message ID: 14300     Entry time: Fri Nov 16 10:53:07 2018     In reply to: 14297     Reply to this: 14302
Author: aaron 
Type: Update 
Category: IOO 
Subject: IMC problematic 

Back to loss measurements.

I replaced the PD I've been using for the AS beam.

I misaligned the x arm.

I tried to lock the y arm, but PRC was locked so I could was unable. Gautam reminded me where the config scripts are.

The armloss measurement script needed two additional modifications:

  • It was setting the initial offset of the PIT and YAW demod signals to 0, but due to the clipping on the heater we are operating at an offset. I commented out these lines.
  • When the script ran UNFREEZE_DITHER, it was running it using medmrun. The scope script hadn't been using this, and it seemed that when it ran UNFREEZE_DITHER in this way the YARM_ASS servo was passing only '0'. I don't really know why this was, but when I removed the call to medmrun it worked.

I ran successfully the loss measurement script for the x and y arms. I'm getting losses of ~100ppm from the first estimates.

I made the following changes to the lossmap script:

  • make the averaging time an input to the script, so we can exceed 2 second averages
  • remove anything about getting data from the scope, replace it with the correct analogues to save the averages for POX/POY refl, MC trans, op lev P/Y, and ASDC signal.
  • record the GPS time in the file with the cds averages (this way I can grab the full data)
  • Added a step in the lossmap script to misalign the optic, so we can continue getting data for the 'misaligned' state, both for the centered and not-centered measurements (that is, for every position on the lossmap).

When the optic aligns itself not at the ideal position, I'm noticing that it often locks on a 01. When the cavity is then misaligned and restored, it can no longer obtain lock. To fix this, I've moved my 'save' commands to just before the loop begins. This means the script may take longer to run, but as long as the cavity is initially locked and well aligned, this should make it more robust against wandering off and never reacquiring lock.

I left the lossmap script running for the x-arm. Next would be to run it for the y arm, but I see that after stepping to a few positions the lock is again lost. It's still trying to run, but if you want to stop it no data already taken will be lost. To stop it, go to the remaining terminal open on rossa and ctrl+c

the analysis needs:

  • Windowing
  • Filter, don't average
  • detrend to get rid of the linear drifts in lock that we see.
    • Is this the right thing?
Attachment 1: Screenshot_from_2018-11-16_19-22-34.png  39 kB  Uploaded Fri Nov 16 19:22:59 2018  | Hide | Hide all
Screenshot_from_2018-11-16_19-22-34.png
ELOG V3.1.3-